1. Compounds containing sulfur and oxygen are serious air pollutants; they represent the major cause of acid rain. Analysis of a sample of a pure compound reveals that it contains 50.1% sulfur and 49.9% oxygen by mass. What is the empiral formula of the compound?

2. A compound with the empirical formula CH2O has a weight of 180 g/mol. What is the molecular formula?

3. What is the molecular formula of a compound that has a molecular weight of 92 and an empirical formula of NO2?
1
Expert's answer
2019-11-11T07:48:11-0500

1. SO2, because M(S) = 2*M(O)

2. C6H12O6, M(CH2O) = 12 + 2 + 16 = 30, Molecular formula is (CH2O)n, where n= 180/30 = 6

3. N2O4, M(NO2) = 14 + 32 = 46, Molecular formula is (NO2)n, where n= 92/46 = 2